OpenStudy (anonymous):

Which of the following equations has no solution? 3x + 1 = 2x + 1 3x + 1 = 3x + 1 3x + 1 = 3x + 2

OpenStudy (anonymous):

3x+1=3x+2 has no solution

OpenStudy (anonymous):

-.-

OpenStudy (anonymous):

what about The product of 8 and the sum of 4 and a number is 112. What is the number? Which equation could be used to solve for the number? 8 + 4n = 112 8 x 4 + n = 112 8(4 + n) = 112

OpenStudy (anonymous):

please medal someone

OpenStudy (anonymous):

8(4+n)=112

OpenStudy (anonymous):

gtg bye and thanks for the medal